Whats Hot G HP Games Imported From IE i Mr. Espings Math Sit S Post-Oka Kinda ease follow the instructions explicitly to get full credit for the answer. se any of the materials from this course, the readings provided, or the HBS case materials. You m quire arch engines or any websites derived therefrom. xpected that you will work on this assignment by yourself. nment is worth 100 points and is due before class on January 30th. Question 15 10 pt If the growth (g) of a company is 2% and the ROIC is 10%, what is the investment rate (IR)? Enter your answer rounded to the nearest whole number, i.e. fty percent would be entered not 0.50. DO NOT include a percentage sign in your answer. Previous Next Quiz saved at 8:13pm Submit Qu
Answer #1

ROIC= Growth/ Investment rate

0.1= 0.02/ IR

IR = 0.02/ 0.1 = 0.2

IR = 20 percent

Answer is 20